katherine:

Write the equation of the line that passes through (3, 4) and (2, −1) in slope-intercept form. y = 3x − 7 y = 3x − 5 y = 5x − 11 y = 5x − 9

katherine:

@SourMunchkin7806

SourMunchkin7806:

i dont know this one sorry lol...i know bits and pieces of this section of math

katherine:

its ok:P

katherine:

@JustSaiyan @Felicity96 @Vocaloid

Felicity96:

Katherine what do you think

katherine:

idk. that's why I'm asking

Felicity96:

Take a guess

katherine:

A

563blackghost:

You need to first identify the slope, plug the points in here... \(\Large\bf{\frac{-1-4}{2-3}}\) what is the slope?

katherine:

umm

katherine:

-5/-1

563blackghost:

Correct, so the negatives cancel to 5. So that is our slope! \(\large\bf{y=5x+b}\) Now we need to identify the y-intercept. So we would plug in one coordinate into the equation. Let's use (3,4). \(\large\bf{4=5(3)+b}\) What is b?

katherine:

y=3x-5

563blackghost:

not quite. First we multiply. \(\large\bf{4=15+b}\) Then subtract 15 from each side. \(\large\bf{-11=b}\) \(\large\bf{y=5x-11}\)

katherine:

So, C?

563blackghost:

yup